OpenStudy (anonymous):

Solve . 20 (-5w-4/20) is greater than or = to -2w

OpenStudy (batman31):

w is less than or equal to -2/49. are you sure you wrote the question right?

OpenStudy (anonymous):

well this is the original equation ... \[20\left[\begin{matrix}1 & w \\ 5 & 4\end{matrix}\right]\ge-2w\]

OpenStudy (anonymous):

you subtract the fractions in the parenthesis .

OpenStudy (anonymous):

@BATMAN31

OpenStudy (anonymous):

@iGreen @chosenmatt